### v1

# Improving Quotation Attribution with Fictional Character Embeddings

Abstract
Humans naturally attribute utterances of direct speech to their speaker in literary works. When attributing quotes, we process contextual information but also access mental representations of characters that we build and revise throughout the narrative. Recent methods to automatically attribute such utterances have explored simulating human logic with deterministic rules or learning new implicit rules with neural networks when processing contextual information. However, these systems inherently lack character representations, which often leads to errors on more challenging examples of attribution: anaphoric and implicit quotes. In this work, we propose to augment a popular quotation attribution system, BookNLP, with character embeddings that encode global information of characters. To build these embeddings, we create DramaCV, a corpus of English drama plays from the 15th to 20th century focused on Character Verification (CV), a task similar to Authorship Verification (AV), that aims at analyzing fictional characters. We train a model similar to the recently proposed AV model, Universal Authorship Representation (UAR), on this dataset, showing that it outperforms concurrent methods of characters embeddings on the CV task and generalizes better to literary novels. Then, through an extensive evaluation on 22 novels, we show that combining BookNLP’s contextual information with our proposed global character embeddings improves the identification of speakers for anaphoric and implicit quotes, reaching state-of-the-art performance. Code and data will be made publicly available.
{NoHyper}
Improving Quotation Attribution with Fictional Character Embeddings
1 Introduction
Authors of literary works employ various devices to create engaging narratives, often combining narration with character dialogues to unveil the plot. Fictional characters portray themselves through dialogues, revealing aspects of their personality, their own style and ideas about themselves and the fictional world. As part of studying characters in digital humanities, automatically identifying utterances and attributing them to characters, also known as quotation attribution, is central Elson et al. (2010); Muzny et al. (2017a); Labatut and Bost (2019); Sims and Bamman (2020).
Recent approaches to quotation attribution typically succeed at attributing explicit quotes but struggle when it comes to anaphoric and implicit quotes Muzny et al. (2017b); Cuesta-Lazaro et al. (2022); Vishnubhotla et al. (2023); Su et al. (2023). Explicit utterances occur when the narrator indicates the speaker of a quote with a speech verb and a named mention, while anaphoric quotes are introduced with a speech verb and a pronoun or common noun. When no narrative information is given on the speaker of the quote, we refer to those as implicit quotes. Examples of such quotes are given in Figure 1.
  1. "As soon as ever Mr. Bingley comes, my dear," said Mrs. Bennet, "you will wait on him of course."
  2. "No, no. You forced me into visiting him last year, and promised if I went to see him, he should marry one of my daughters…"
  3. His wife represented to him how absolutely necessary such an attention would be from all the neighbouring gentlemen, on his returning to Netherfield.
  4. "’Tis an etiquette I despise," said he.
Figure 1: Excerpt of Pride and Prejudice by Jane Austen (1813). Quotations are colored by quote type: explict, implicit and anaphoric. Speaker information given by the narrator are underlined.
Attributing anaphoric and implicit quotes usually requires fine-grained understanding of contextual information such as linking pronouns to their canonical entity or discerning discussion patterns. Accessing information of fictional characters (style, persona, gender) also help disambiguate challenging examples of attribution. Variations of writing style and long-range contextual dependencies, ubiquitous in fictional works, add an extra layer of complexity, making attribution of non-explicit quotes a complex task to solve even with recent tools. Sims and Bamman (2020); Vishnubhotla et al. (2022)
While some of these shortcomings have been addressed by improving link prediction between pronouns/common nouns and canonical characters Vishnubhotla et al. (2023) or by exploiting better contextual representations Cuesta-Lazaro et al. (2022); Su et al. (2023), we instead show that combining representations encoding global stylistic information and topical preferences of fictional characters with contextual information improves the identification of speakers of anaphoric and implicit quotes. In particular, we combine these embeddings with BookNLP [^1] https://github.com/booknlp/booknlp, a widely used NLP pipeline designed for literary texts. As explicit utterances can easily be attributed to their speaker, we assume those quotes given, and use them to inject global character-level information into BookNLP.
Among methods to represent characters from the words they utter, previous works Li et al. (2023); Aggazzotti et al. (2023); Michel et al. (2024) have explored embeddings encoding topical preferences using SBERT Reimers and Gurevych (2019) or stylistic information using Universal Authorship Representations (UAR) Rivera-Soto et al. (2021), a fine-tuned variant of SBERT trained on the Authorship Verification (AV) task with millions of Reddit users. However, considering the discrepancies between Reddit and literary discourse, we question the effectiveness of these embeddings to represent characters. To overcome this potential limitation, we build a variant of UAR tailored for literary texts and evaluate those representation models on quotation attribution and Character Verification (CV), a task similar to AV, aiming at distinguishing characters in literary works using their utterances within authorial bounds.
Our proposed models significantly improve CV on drama plays and novels, and reaches state-of-the-art quotation attribution accuracy of non-explicit quotes when used in combination with BookNLP. Besides, we show that using models tailored for literary text leads to better quotation attribution performance. To sum up, our contributions are:
  • We construct DramaCV, a corpus of English drama plays, dating from the 15th to 20th century focused on CV, and train variants of UAR on this dataset, showing that our proposed models significantly improves CV on DramaCV and also generalize better to novels.
  • Through an extensive evaluation, we show that BookNLP can be easily extended to improve its current performance on the Project Dialogism Novel Corpus (PDNC) Vishnubhotla et al. (2022), and reaches state-of-the-art accuracy on non-explicit quotes when trained in combination with global character representations computed from their explicit quotes.
2 Related Work
Fictional Character Embeddings.   Understanding fictional characters objectives, personas and style often comes naturally when reading novels. When accessing information of the fictional world and interactions between characters provided throughout the narrative, we infer mental representations of characters that are revised when accessing new information Gernsbacher et al. (1992); Culpeper (1996). Previous computational works on this aspect of fictional character understanding have explored representing fictional characters in a latent space, often targeting a specific aspect, such as persona, style or descriptive elements. These computational representations of characters are often used for further analysis on large literary corpora, trying to understand character distinctiveness between authors Bamman et al. (2014) or narrative comprehension Brahman et al. (2021). Other works focus on using fictional characters embedings to improve character-level tasks such as coreference resolution and character linking Li et al. (2023) or contextual and book-level tasks such as narrative question answering Inoue et al. (2022).
Here, we propose to represent character from what characters say and how they say it, building literary embeddings encoding style and topical preferences of characters, derived from models trained on Authorship Verification or Character Verification. We define Character Verification (CV) as a task closely related to Authorship Verification, aiming at distinguishing characters in literary works, within authorial bounds. Most works focusing on CV analyze drama characters because of the availability of large annotated corpora and aim at capturing syntactic, lexical and phonological variations that occur in their direct speech using statistical measures of distinctiveness Dinu and Uban (2017); Vishnubhotla et al. (2019); Šeļa et al. (2023).
With the recent introduction of PDNC Vishnubhotla et al. (2022), a corpus of 28 novels annotated with quotation attribution, analysis on fictional characters in novels can be done at a larger scale. Michel et al. (2024) provide a comparative study of PDNC’s characters, converting the original PDNC quotation attribution task to a CV task and showing that UAR Rivera-Soto et al. (2021) performs well on the task. However, this model is trained on Reddit posts, which are inherently different from fictional character utterances, motivating the need for an in-domain variant of UAR.
text
Quotation Attribution.   Datasets of quotation attribution containing fully annotated English novels are rather scarce, and are often limited to one or two books. Elson and McKeown (2010) introduce the CQSC corpus and attribute automatically extracted quotes to named entities and nominals with a supervised mention ranking model. Instead, He et al. (2013) attribute quotes directly to speakers with a supervised ranking system using features such as speaker alternation patterns and character-level features He et al. (2010). Other approaches involve sequence labeling O’Keefe et al. (2012) and Dialogue State Tracking Cuesta-Lazaro et al. (2022). The deterministic sieve-based model of Muzny et al. (2017b) regards quotation attribution as a two-step process: quote-mention linking and mention-speaker linking. The NLP pipeline dedicated to books, BookNLP, replaced the deterministic sieves with fine-tuned language models. The link between mention and speaker is carried out by BookNLP’s coreference resolution model, and the quote-mention linking is predicted by a BERT model, fine-tuned on LitBank’s speaker annotations Sims and Bamman (2020). Recently, Vishnubhotla et al. (2022) introduce the largest-to-date corpus of quotation attribution, PDNC, and show a similar accuracy score of around 63% for both BookNLP and the sieve-based model. However, better results were obtained later by fine-tuning BookNLP on PDNC and restricting the predicted coreference chains to resolved characters only Vishnubhotla et al. (2023). SIG Su et al. (2023) adapted quotation attribution for encoder-decoder models, showing that predicting both the speaker and the addressees of a quote with BART reaches state-of-the-art accuracy on PDNC.
Interestingly, reported BookNLP attribution results on PDNC often approach perfect accuracy for explicit quotes, indicating that most of the research efforts should be put in the attribution of non-explicit quotes. Driven by this insight, we exploit explicit quotes uttered by characters to represent them in a latent space, and are the first to show their potential for quotation attribution when combined with contextual representations.
3 Literary Character Representations
In this section, we introduce DramaCV (Drama Character Verification), a dataset of English drama plays focused on CV. Then, we compare variants of Universal Authorship Representation (UAR) trained on DramaCV with the original UAR and SBERT. We finish by analyzing domain adaptation results on the CV variant of PDNC.
Split Segments Utterances Queries Targets/Query (avg)
Train 1507 263270 5392 5.0
Scene Val 240 50670 1557 8.8
Test 203 41830 1319 8.7
Train 226 449407 4109 90.7
Play Val 30 63934 917 55.1
Test 31 74738 1214 108.5
PDNC Explicit - 6303 562 11.2
Table 1: Summary statistics of the Scene and Play instances of DramaCV by split and the Explicit split of Michel et al. (2024) that we use to test generalization.
Among models trained on Authorship Verification, UAR, a fine-tuned variant of SBERT, is able to distinguish millions of Reddit authors. It learns a representation space under which a collection of documents composed by an author has a higher similarity with another collection from the same author than with a collection composed by another author Rivera-Soto et al. (2021). Although the Reddit version of UAR appears to perform well on PDNC, which is on-par with its zero-shot domain transfer abilities Rivera-Soto et al. (2021), we question its suitability for literary analysis and propose to train instead a domain-specific variant of UAR tailored for literary texts. Reddit has the advantage of spanning millions of different authors, but posts are often written informally and exhibit strong stylistic cues such as typographical or case errors Wegmann et al. (2022). In contrast, literary texts do not contain such errors and often span a wider vocabulary that evolves over time, in concordance with writing styles. This motivates the need for an in-domain UAR for analyzing authorship of fictional characters. Besides, quotes are subject to the writing style of authors. We would like character representations disentangled from the author’s stylistic preferences.
3.1 DramaCV
Character Verification aims at answering the following problem: given two collections of quotes that have been uttered by characters within the same work, predict whether they were spoken by the same character or not. Although closely linked to Authorship Attribution, we note that this framework is different for the following reasons: 1) we do not have access to a list of canonical characters for every literary work and 2) we compare characters within the same story, implying that each quote is inherently influenced by the author’s style. In comparison, Authorship Attribution aims at determining who wrote a text among a list of predefined authors, which is also inherently different from quotation attribution. Multiple documents come naturally annotated with who wrote or spoke them, such as social media Baumgartner et al. (2020) movie scripts Sang et al. (2022) or drama plays Fischer et al. (2019). Among existing datasets, DraCor Fischer et al. (2019) contains drama plays in various languages, but only provides Shakespeare plays for English language. Alternatively, explicit utterances from PDNC are an additional data source, but we wanted to evaluate the generalization of literary character embeddings for quotation attribution to unseen novels. Hence, we instead use the Explicit split of PDNC Michel et al. (2024) as an additional test split to analyze generalization from drama to novels. To the best of our knowledge, no datasets focusing on English drama and containing speaker annotation has been publicly released.
We create DramaCV, a publicly available dataset of drama plays written by English authors, dating from the 15th to 20th century. We extract 499 plays from Project Gutenberg [^2] https://gutenberg.org/ with the GutenTag software Brooke et al. (2015) that we parsed to attribute each line to its speaker. Then, we construct two instances of DramaCV:
Scene: We split each plays in scenes, a small segment unit of drama that is supposed to contain actions occurring at a specific time and place with the same characters. If a play has no <scene> tag, we instead split it in acts, with the <act> tag. Acts are larger segment units, composed of multiple scenes. We do not consider plays that have neither of these tags. Although scenes usually contain a small number of characters, they represent a challenge for CV models as participating characters are akin to discuss the same topic. We expect a model trained on scenes to rely mostly on stylistic cues.
Play: Here, we do not segment plays, and instead use all character lines in the full play. Compared to the scene instance, all characters participating in the play are considered as targets for the AV task. Here, topical preferences of each character across the whole play can be used in addition to style.
We parsed a total of 169 plays for Scene, and 287 plays for Play. The resulting plays are split in 80% for training, 10% for validation and 10% for test. For validation and test, half utterances of a character are randomly sampled for the query, and the remaining half for the target. For the train split, we consider as queries and targets 8 utterances from the pool of a character’s set of lines, that we randomly sample at each epoch. We provide summary statistics for both instances of DramaCV in Table 1.
3.2 Experimental Setup
We train UAR from scratch using its public implementation on DramaCV. For both instances, we use the pre-trained all-distilroberta-v1 as the base encoder, and employ a learning rate of 2e-5 for 20 epochs. We set the dimension of the model to $d=512$ and use a maximum sequence length of 64 tokens for all utterances. During training, we build queries and targets with distinct collections of 8 randomly sampled utterances to represent characters. A batch size of 1 play is used when training on the Play split, and 8 scenes are used when training on the Scene split. During inference, we do not restrict the number of utterances in a collection.
A supervised contrastive objective Khosla et al. (2020) is used where in-batch negatives considered are characters within the same segment unit. We motivate this choice for each dataset instance:
Scene.  In a scene, characters are likely to discuss similar topics. By restricting the objective to characters within a scene, we force the model to rely on stylistic cues to distinguish lines of characters.
Play. We want to avoid capturing the writing style of authors to distinguish among collections of character lines. Thus, we restrict the set of in-batch negatives to characters within the same work.
We compare our variants of UAR, $\text{UAR}_{Scene}$ and $\text{UAR}_{Play}$ against their Reddit counterpart, $\text{UAR}_{Reddit}$ and SBERT. Following previous work Aggazzotti et al. (2023); Michel et al. (2024), we create SBERT character embeddings by encoding each utterance in a vector with a pretrained model [^3] We use all-mpnet-base-v2., followed by a coordinate-wise mean of all utterance vectors. Instead, UAR directly encodes a collection of utterances into a $d$-dimensional vector.
For both test-splits of DramaCV, we encode a query with one of the mentioned models, yielding a character representation, and then encode each associated targets, which are collections of lines uttered by all characters in the same segment unit (including other lines from the same character). We then compute the cosine similarities between a query and the targets, and evaluate the ranking performance of these embeddings with the Area Under the Receiver Operating Characteristic curve (AUC) metric. We also test generalization to novels with the Explicit split of PDNC. In this experimental setup, Michel et al. (2024) proposed two evaluations: Character-Character (CC) and Character-Quote (CQ). The task in CC is to answer the following problem: given a collection of all explicit quotes of a character in a chapter, and a collection of quotes of any type from other chapters, determine if they have been spoken by the same character or not. The CQ experiment is constructed with the same collections of explicit quotes, but we instead need to predict if a single quote of any type from other chapters has been spoken by the same character or not. The later experiment resembles quotation attribution, but no contextual information is provided.
DramaCV PDNC - Exp
Scenes Plays CC CQ
SBERT 72.0 (5) 78.8 (10) 63.8 (16) 54.4 (5)
$\text{UAR}_{Reddit}$ 68.6 (7) 74.5 (11) 81.1 (9) 54.1 (4)
$\text{UAR}_{Scene}$ 82.3 (6) - 81.3 (10) 58.9 (4)
$\text{UAR}_{Play}$ 84.2 (9) 86.5 (8) 58.5 (5)
Table 2: Average AUC (%) on test plays in DramaCV and novels in the Explicit split of PDNC with standard deviations between parentheses.
3.3 Results
We report AUC for both test splits of DramaCV and the Explicit split of PDNC in Table 2. Interestingly, $\text{UAR}_{Reddit}$ performs worse on both splits compared to SBERT and in-domain UARs. $\text{UAR}_{Scene}$ increases AUC on its training domain up to 10 points compared to SBERT, indicating that the model is able to capture stylistic cues specific to drama that help distinguishing utterances of characters discussing similar topics. The increase of $\text{UAR}_{Play}$ is slightly lower, but the high AUC indicates that it succeeds at distinguishing characters within the same play. The high standard deviation across plays suggests that task difficulty is not distributed uniformly among these literary works. Among best-performing plays, we found works by canonical authors such as Shakespeare’s Hamlet, Shaw’s John Bull’s Other Island or Ben Jonson’s The Alchemist (see complete list in Appendix A).
Looking at generalization results to PDNC, $\text{UAR}_{Reddit}$ performs better than SBERT, but almost equals $\text{UAR}_{Scene}$ in the CC setup. We find significant improvement with $\text{UAR}_{Play}$ over the baselines, suggesting that training UAR on drama allows better domain adaptation to literary novels. We also remark that this adaptation requires less data, as training $\text{UAR}_{Scene}$ with around 5000 drama characters is enough to match the performance of $\text{UAR}_{Reddit}$ that has been trained with millions of different users.
The advantages of domain adaptation can also be seen in the CQ evaluation setup, where the two variants of UAR trained on drama show a significant increase in AUC over the baselines. This indicates that character representations built with our proposed UARs are more similar to embeddings of their other quotes than representations built with $\text{UAR}_{Reddit}$ and SBERT. However, AUC scores for the CQ experiment are low, showing that the task of attributing quotes requires more than only comparing similarities of character and quote embeddings.
Motivated by the ability of UAR to encode character utterances in a representation space that contain useful, though insufficient information for quotation attribution, we evaluate next the impact of combining these global character embeddings with BookNLP’s contextual information.
4 Quotation attribution
Non-Explicit Explicit Anaphoric Implicit
SIG $70.0^{+}$
ChatGPT $70.0^{+}$
Unanswerable (%) 20.9 (6.7) 0 24.5 (8.1) 19.8 (8.6)
[…]
+ SBERT $70.0^{\star\phantom{\star}}$ (4.5) 98.3 (1.1) $70.8^{\phantom{\star\star}}$ (7.3) $68.1^{\star}$ (5.5)
+ $\text{UAR}_{Reddit}$ $71.1^{\star\star}$ (4.7) 98.7 (1.1) $71.5^{\phantom{\star\star}}$ (7.9) $\mathbf{69.6}^{\star}$ (5.3)
+ $\text{UAR}_{Play}$ $70.9^{\star\phantom{\star}}$ (4.4) 98.7 (1.0) $71.1^{\phantom{\star\star}}$ (7.9) $69.5^{\star}$ (5.7)
+ $\text{UAR}_{Scene}$ $\mathbf{71.2}^{\star\phantom{\star}}$ (4.3) 98.7 (1.0) $\mathbf{71.7}^{\star\star}$ (8.1) $\mathbf{69.6}^{\star}$ (5.4)
[…]
In this section, we first review BookNLP’s approach to quotation attribution. BookNLP is a popular Natural Language Processing pipeline dedicated to books. Then, we present our straightforward approach to combine BookNLP’s contextual information with embeddings of characters computed from their explicit quotes and finish by discussing our main results.
4.1 BookNLP
BookNLP’s pipeline includes among others Part-of-speech tagging, Named Entity Recognition, Coreference Resolution and quotation attribution. All components of the pipeline have been trained on LitBank Bamman et al. (2019, 2020); Sims and Bamman (2020), a corpus containing 100 chapters from public domain books. Since its quotation attribution model is not part of a publication, we will review its details in the following paragraphs.
BookNLP carries out quotation attribution in four independents steps: 1) tokenization and named entity recognition, 2) character name clustering, 3) pronominal coreference resolution, and 4) quote-mention linking. The pronominal coreference resolution step outputs mentions along with the predicted character they refer to. Using predicted mentions in the surrounding context of a quote, the quote-mention linking step aims at predicting which mention is the one that refers to the true speaker of the quote. It is achieved by a fine-tuned BERT model Devlin et al. (2019). We give further details of the quote-mention linking step below.
Let $D=(t_{1},\dots t_{n})$ be a tokenized document and $q=(t_{i},\dots,t_{j})$ be a character utterance in this document, starting at token $i$ and ending at token $j$. A contextual segment is computed for each quote, $c_{q}=[\,c_{q}^{left}\;|\,q\,|\;c_{q}^{right}\,]$, where $c_{q}^{left}$ and $c_{q}^{right}$ are left and right contextual information of length $w$ tokens and [|][\,|\,][ | ] denotes concatenation. Additionally, the full quote $q$ is replaced by a special token [QUOTE] and all other quotes occurring in the contextual segment are replaced by the special token [ALTQUOTE], such that the contextual segment mainly reflect narration. Each segment $c_{q}$, containing masked quotes and narrative parts, is fed to BERT, producing a contextual representation $\mathbf{H}_{c_{q}}$. Then an unary compatibility score between the (masked) quote token $q$ and a mention-span $m=(t_{m_{s}},\dots,t_{m_{e}})$ occurring in the segment is calculated as follows:
$h_{m}$ $=\frac{1}{|m|}\sum\limits_{i=m_{s}}^{m_{e}}\mathbf{H}_{c_{q}}[i]$ (1)
$s(q,m)$ $=\phi([\mathbf{H}_{c_{q}}[q]\;|\;h_{m}])$ (2)
where $\mathbf{H}_{c_{q}}[i]$ is the BERT representation of token $i$ while $|m|$ is the length in tokens of mention $m$ and $\phi$ is a feed-forward neural network with one hidden layer and a tanh activation. Given a set of candidates $Y(q)=(m_{1},\dots,m_{k})$, the model is trained to optimize the log-likelihood of mentions referring to the true speaker of $q$.
The default window size is $w=50$ tokens, and the hidden layer dimension is $100$. During training and inference, only a maximum of $k=10$ mention candidates are considered. In our experiments, we modify these default parameters by increasing the narrative window size to $w=100$ tokens, and using a relu activation and a hidden dimension of $d=512$ in the feed-forward neural network. We also replace BERT by SpanBERT-large Joshi et al. (2020), computing mention representations by concatenating SpanBERT embeddings of the first and last tokens. At inference, we do not truncate the number of candidates and instead use all candidate mentions available in the context window.
4.2 Integrating Character Embeddings
Let $E(c)=(q_{i},\dots,q_{j})$ be the collection of all explicit quotes uttered by character $c$. We embed this collection of quotes using one of the model described in Section 3.2, yielding a vector representation $v_{c}$ for each character. If a character has not been explicitly quoted, we set its vector representation to zeros. Additionally, we embed each quote $q$ independently with the same model to get a vector representation $u_{q}$. We modify Equation 2 to combine BookNLP’s contextual representation with these character and quote representations:
$s(q,m)=\phi([\mathbf{H}_{c_{q}}[q]\;|\;h_{m}\;|\;v_{\text{entity}(m)}\;|\;u_{q}])$ (3)
where $\text{entity}(m)$ is the character predicted by the coreference resolution step for mention $m$.
Equation 3 allows BookNLP to rely on extra-contextual information by injecting global character information and compatibility between quote $q$ and candidate character $c$ derived from CV models. In this work, our primary goal is to prove the usefulness of character embeddings. We leave as future work finding the best strategy to integrate them in systems such as BookNLP.
4.3 Experimental Setup
Following Su et al. (2023), we use the first version of PDNC, which contains 22 annotated novels. We focus on major and intermediate characters who are characters that uttered at least 10 quotes in the novel. The later allows to avoid the long tail of minor characters and to compare our results to the current state-of-the-art, SIG and ChatGPT.
We base our implementation on BookNLP+ Vishnubhotla et al. (2023), that restricts coreference predictions to chains that can be successfully resolved to a character entity from the annotated PDNC character lists. We follow the same evaluation protocol, using their provided cross-validation split at the novel level, reporting the average accuracy over each split.
[…]
4.4 Results
Our main results are displayed in Table 3. Our implementation of BookNLP+ reaches 68.9% accuracy on non-explicit quotes, compared to the 53% reported in Vishnubhotla et al. (2023). Combining BookNLP+ with SBERT character embeddings matches the accuracy of SIG and ChatGPT on non-explicit quotes. The gains are larger for all variants of UAR, reaching 71.2% accuracy on non-explicit quotes with our proposed $\text{UAR}_{Scene}$. As expected, we found that these gains are largely due to the improved attribution performance of implicit quotes, increasing the accuracy by more than 3 points with all UAR models.
$\Delta$ Non-Exp $\Delta$ Ana $\Delta$ Imp
SBERT $-0.0\phantom{{}^{\star\star}}$ $-0.0\phantom{{}^{\star}}$ $-0.1\phantom{{}^{\star}}$
$\text{UAR}_{Reddit}$ $-0.1\phantom{{}^{\star\star}}$ $-0.0\phantom{{}^{\star}}$ $-0.2\phantom{{}^{\star}}$
$\text{UAR}_{Play}$ $-0.2^{\star\star}$ $-0.1\phantom{{}^{\star}}$ $-0.3^{\star}$
$\text{UAR}_{Scene}$ $+0.1\phantom{{}^{\star\star}}$ $+0.1\phantom{{}^{\star}}$ $+0.1\phantom{{}^{\star}}$
Table 4: Variation of accuracy when using predicted speaker of explicit quotes. Statistical significance from paired t-test is indicated with ${}^{\star}=5\%$, ${}^{\star\star}=10\%$
Interestingly, although our proposed variants of UAR tailored for literary texts are better in the CV task, we only see a small – though statistically significant – increase between $\text{UAR}_{Scene}$ and its Reddit counterpart on non-explicit attribution accuracy. The reasons could be twofold. First, BookNLP’s system only focuses on finding the right narrative information that helps understanding the speaker of a quote. However, our embeddings mostly reveal stylistic cues of characters, which are not encoded by BookNLP. Therefore, our implementation provide external character information that might not align perfectly with BookNLP’s contextual decisions, limiting their combination potential. Second, we add character information based on noisy coreference chains, which might hinder the model’s peformance. We believe further research should be conducted in finding better combinations between character embeddings and BookNLP.
The high standard deviations across novels reveals that the task is harder for some novels than others. Looking at the percentage of unanswerable quotes (quotes that do not have a candidate mention referring to their speaker in context) provides a reason for such high variations: it ranges from 0 to 56% per novel. Figure 2 shows the attribution accuracy on non-explicit quotes when only considering novels that have a low percentage of unanswerable quotes. Only 4 novels have a percentage lower than 5%, but 17 out of the 22 have a percentage lower than 20%. As expected, we see higher performance for all models, and the gain from UAR character embeddings is more clear.
Non-Exp Ana Imp
Unanswerable (%) 20 21 18
[…]
+ SBERT $69.3^{\phantom{\star\star}}$ $71.6^{\phantom{\star\star}}$ $68.0^{\phantom{\star}}$
+ $\text{UAR}_{Reddit}$ $70.5^{\star\star}$ $72.4^{\phantom{\star\star}}$ $69.4^{\star}$
+ $\text{UAR}_{Play}$ $71.1^{\star\phantom{\star}}$ $72.8^{\star\star}$ $\mathbf{70.1}^{\star}$
+ $\text{UAR}_{Scene}$ $\mathbf{71.2}^{\star\phantom{\star}}$ $\mathbf{73.1}^{\phantom{\star\star}}$ $70.0^{\star}$
[…]
Let us note that $\text{UAR}_{Scene}$ is consistently better than $\text{UAR}_{Play}$, though the latter is better at distinguishing PDNC characters in the CC experiment. Restricting UAR to encode stylistic information by using collections of quotes with similar topic seems to generalize better to quotation attribution.
[…]
Finally, we present quotation attribution results on 6 annotated novels from the second release of PDNC in Table 5. We use the model trained on the first cross-validation split to predict attributions on each of these novels. The impact of using literary character embeddings is more clear: both of our proposed variants of UAR show increased attribution accuracy against their Reddit counterpart.
Figure 2: Accuracy results on non-explicit quotes averaged over subsets of novels that have a low percentage of unanswerable quotes, with standard deviations.
5 Conclusion
In this work, we evaluated fictional character embeddings constructed with the explicit quotes they utter, showing that they improve quotation attribution accuracy of non-explicit quotes on PDNC when used in combination with BookNLP’s contextual representations. We created DramaCV, a corpus of English drama plays focused on Character Verification and trained two variants of UAR on this dataset. Our proposed models show better performance when verifying fictional characters on both DramaCV and PDNC. Character embeddings derived from these models are further used for quotation attribution on PDNC, reaching state-of-the-art accuracy when combined with BookNLP+. The impact of using literary character embeddings against embeddings derived from a model trained on Reddit is small, though statistically significant, suggesting superiority of in-domain character representations for quotation attribution.
We proposed a straightforward way to combine fictional character embeddings with contextual representations, using BookNLP as an example. Our approach can be adapted to other models: cross-attention over character embeddings could be used for encoder-decoder models, while modality alignment Zhang et al. (2023) is a solution for decoder-only models. We leave potential improvements to the combination of fictional character embeddings and contextual representations to future work.
We found that BookNLP’s attribution accuracy is not distributed uniformly across novels, mainly caused by noisy coreference resolution resulting in many quotes not having candidate mentions that refer to their true speaker. We believe that including our character embeddings in the coreference resolution model as in Li et al. (2023) might improve its performance and thus increase overall attribution accuracy. Another potential solution could include unmasking the closest quotes in the vicinity of a target quote that are likely to contain addressee information. We leave better uses of fictional character embeddings with BookNLP for future exploration.
6 Limitations
In this work, we focused on integrating fictional character embeddings in encoder-only models. We have not tested with other types of models such as encoder-decoder and decoder only, but we believe that integrating our embeddings in those framework is possible and is likely to improve their performance on the quotation attribution task. Our first set of experiments corroborate this assumption, on the usefulness of our proposed character embeddings in distinguishing fictional characters.
We also proposed models that can be used to create character representations. These models are trained on a corpus of drama plays, which spans almost 400 years of works. It is unclear how the variation of writing style that have occurred throughout these centuries affects the information contained in these embeddings, and how they can generalize to new literary styles. This remark also extends to our results on quotation attribution. We presented results on a corpus containing novels from the 19th to early 20th century ; which is closer in terms of writing style to DramaCV than newer novels published in the 21th century. It is thus unclear how our results generalize to these newer novels, but we strongly believe that accurate character embeddings improve quotation attribution regardless of the writing style of the author.
We only focused on representations encoding stylistic cues and topical preferences and showed that they improve quotation attribution. However, this type of character embeddings might be incomplete and other aspects of characters can be used in addition to our proposed embeddings (personas, gender, objective, or social status). This might allow to build accurate representations of characters, that are likely to improve even more quotation attribution accuracy.
Appendix A Performance on all Test Plays.
We report performance of $\text{UAR}_{Scene}$ and $\text{UAR}_{Play}$ on their respective test split in Table 6 and Table 7.
Appendix B Computing Information
[…]
Play Author AUC
Justice John Galsworthy 95.6
The Alchemist Ben Jonson 91.8
Hamlet William Shakespeare 86.5
Plays : Third Series John Galsworthy 86.3
The works of John Dryden John Dryden 85.2
The Life of Timon of Athens William Shakespeare 84.2
King Henry the Fifth William Shakespeare 83.7
Night Must Fall : a Play in Three Acts Emlyn Williams 83.7
The Spanish Tragedy Thomas Kyd 83.2
The works of John Dryden John Dryden 83.0
The Scarlet Stigma: A Drama in Four Acts Nathaniel Hawthorne 82.4
King John William Shakespeare 81.8
The Works of Charles and Mary Lamb Mary Lamb 80.8
The Scornful Lady John Fletcher 80.8
Married Life A Comedy, in Three Acts John Baldwin Buckstone 80.0
The Flutter of the Goldleaf; and Other Plays Frederick Peterson 77.5
Single Life A Comedy, in Three Acts John Baldwin Buckstone 75.9
The Countess Cathleen W. B. (William Butler) Yeats 75.6
De Turkey and De Law: A Comedy in Three Acts Zora Neale Hurston 66.4
Table 6: AUC per novel on the Scene test split of DramaCV.
Play Author AUC
Bride Roses William Dean Howells 100.0
Fanny’s First Play Bernard Shaw 100.0
Five Little Plays Alfred Sutro 98.6
John Bull’s Other Island Bernard Shaw 95.4
Mrs. Dot: A Farce William Somerset Maugham 94.5
The Earl of Essex: A Tragedy, in Five Acts Henry Jones 94.4
The Way of the World William Congreve 92.7
Plays : Fifth Series John Galsworthy 92.5
Representative Plays by American Dramatists: 1856-1911: The New York Idea Langdon Elwyn Mitchell 91.9
Plays: Lady Frederick, The Explorer, A Man of Honour William Somerset Maugham 90.3
Great Catherine Bernard Shaw 89.3
Hypolympia Or, The Gods in the Island, an Ironic Fantasy Edmund Gosse 87.1
The works of John Dryden John Dryden 87.0
The Grecian Daughter Arthur Murphy 85.5
Dramatic Technique George Pierce Baker 81.6
Lyre and Lancet: A Story in Scenes F. Anstey 80.2
Beaumont and Fletcher’s Works, Vol. 5 Francis Beaumont 79.0
King Henry VI, Part 3 William Shakespeare 78.1
Plays and Lyrics Cale Young Rice 78.0
The Easiest Way Representative Plays by American Dramatists: 1856-1911 Eugene Walter 77.6
Charles Di Tocca: A Tragedy Cale Young Rice 77.3
St. Patrick’s day, or, the scheming lieutenant : a farce in one act Richard Brinsley Sheridan 76.7
Deirdre of the Sorrows John Millington Synge 76.3
Ambrose Gwinett or, a sea-side story : a melo-drama, in three acts Douglas William Jerrold 75.8
Henry V William Shakespeare 75.3
Box and Cox: A Romance of Real Life in One Act. John Maddison Morton 75.0
The Recruiting Officer George Farquhar 74.7
Yolanda of Cyprus Cale Young Rice 74.3
Table 7: AUC per novel on the Play test split of DramaCV.