### v2

# Improving Quotation Attribution with Fictional Character Embeddings

Abstract
Humans naturally attribute utterances of direct speech to their speaker in literary works. When attributing quotes, we process contextual information but also access mental representations of characters that we build and revise throughout the narrative. Recent methods to automatically attribute such utterances have explored simulating human logic with deterministic rules or learning new implicit rules with neural networks when processing contextual information. However, these systems inherently lack character representations, which often leads to errors in more challenging examples of attribution: anaphoric and implicit quotes. In this work, we propose to augment a popular quotation attribution system, BookNLP, with character embeddings that encode global stylistic information of characters derived from an off-the-shelf stylometric model, Universal Authorship Representation (UAR). We create DramaCV [^1] Code and data can be found at https://github.com/deezer/character_embeddings_qa, a corpus of English drama plays from the 15th to 20th century that we automatically annotate for Authorship Verification of fictional characters utterances, and release two versions of UAR trained on DramaCV, that are tailored for literary characters analysis. Then, through an extensive evaluation on 28 novels, we show that combining BookNLP’s contextual information with our proposed global character embeddings improves the identification of speakers for anaphoric and implicit quotes, reaching state-of-the-art performance.
{NoHyper}
Improving Quotation Attribution with Fictional Character Embeddings
1 Introduction
Authors of literary works employ various devices to create engaging narratives, often combining narration with character dialogues to unveil the plot. Fictional characters portray themselves through dialogues, revealing aspects of their personality, their own style and ideas about themselves and the fictional world. As part of studying characters in digital humanities, automatically identifying utterances and attributing them to characters, also known as quotation attribution, is central Elson et al. (2010); Muzny et al. (2017a); Labatut and Bost (2019); Sims and Bamman (2020).
Recent approaches to quotation attribution typically succeed at attributing explicit quotes but struggle when it comes to anaphoric and implicit quotes Muzny et al. (2017b); Cuesta-Lazaro et al. (2022); Vishnubhotla et al. (2023); Su et al. (2023). Explicit utterances occur when the narrator indicates the speaker of a quote with a speech verb and a named mention, while anaphoric quotes are introduced with a speech verb and a pronoun or common noun. When no narrative information is given about the speaker of the quote, we refer to those as implicit quotes. Examples of such quotes are given in Figure 1.
  1. "As soon as ever Mr. Bingley comes, my dear," said Mrs. Bennet, "you will wait on him of course."
  2. "No, no. You forced me into visiting him last year, and promised if I went to see him, he should marry one of my daughters…"
  3. His wife represented to him how absolutely necessary such an attention would be from all the neighbouring gentlemen, on his returning to Netherfield.
  4. "’Tis an etiquette I despise," said he.
Figure 1: Excerpt of Pride and Prejudice by Jane Austen (1813). Quotations are colored by quote type: explict, implicit and anaphoric. Speaker information given by the narrator are underlined.
Attributing anaphoric and implicit quotes usually requires fine-grained understanding of contextual information such as discerning discussion patterns and long-range information such as linking pronouns to their canonical entity. Accessing information of fictional characters (style, persona, gender) also helps disambiguate challenging examples of attribution. Variations of writing style and long-range contextual dependencies, ubiquitous in fictional works, add an extra layer of complexity, making attribution of non-explicit quotes a complex task to solve even with recent tools. Sims and Bamman (2020); Vishnubhotla et al. (2022)
Some of these shortcomings have been addressed by improving link prediction between pronouns/common nouns and canonical characters Vishnubhotla et al. (2023) or by exploiting better contextual representations Cuesta-Lazaro et al. (2022); Su et al. (2023). Instead, we show that combining fictional character representations encoding global stylistic information and topical preferences with contextual information processed by BookNLP,  [^2] https://github.com/booknlp/booknlp, a widely used NLP pipeline designed for literary texts, improves the speaker attribution of anaphoric and implicit quotes. As explicit utterances are often accurately attributed to their speaker Vishnubhotla et al. (2023), we assume those quotes given, and use them to derive stylistic information of characters.
Among methods to represent characters from the words they utter, previous works Li et al. (2023); Aggazzotti et al. (2023); Michel et al. (2024) have explored embeddings encoding topical preferences using SBERT Reimers and Gurevych (2019) or stylistic information using Universal Authorship Representations (UAR) Rivera-Soto et al. (2021), a fine-tuned variant of SBERT trained on the Authorship Verification (AV) task with millions of Reddit users. UAR’s stylometric embeddings built from character utterances were previously shown to distinguish fictional characters among the same work Michel et al. (2024) but considering the discrepancies between Reddit and literary discourse, we question the effectiveness of these embeddings to represent characters. To overcome this potential limitation, we build variants of UAR tailored for literary texts that we validate on the AV task with drama and novel characters.
Stylometric character representations computed with our UAR variants significantly improve AV of characters in literary works, while also reaching state-of-the-art quotation attribution accuracy of non-explicit quotes when combined with BookNLP contextual representations. To sum up, our contributions are:
  • We construct DramaCV, a corpus of English drama plays, dating from the 15th to 20th century, focused on the AV of fictional characters within the same work, and train variants of UAR on this dataset, showing that our proposed models significantly improve AV on DramaCV and also generalize better to novels.
  • Through an extensive evaluation, we show that BookNLP can be easily extended to improve its current performance on the Project Dialogism Novel Corpus (PDNC) Vishnubhotla et al. (2022), and reaches state-of-the-art attribution accuracy of non-explicit quotes when trained in combination with global character representations computed from their explicit quotes.
2 Related Work
Quotation Attribution.
Datasets of quotation attribution containing fully annotated English novels are rather scarce, and are often limited to one or two books. Elson and McKeown (2010) introduce the CQSC corpus and attribute automatically extracted quotes to named entities and nominals with a supervised mention ranking model. Instead, He et al. (2013) attribute quotes directly to speakers with a supervised ranking system using features such as speaker alternation patterns and character-level features He et al. (2010). Other approaches involve sequence labeling O’Keefe et al. (2012), Dialogue State Tracking Cuesta-Lazaro et al. (2022) on English novels, or scoring systems on Chinese novels Chen2021. The deterministic sieve-based model of Muzny et al. (2017b) regards quotation attribution as a two-step process: quote-mention linking and mention-speaker linking. The NLP pipeline dedicated to books, BookNLP, replaced the deterministic sieves with fine-tuned language models. The link between mention and speaker is carried out by BookNLP’s coreference resolution model, and the quote-mention linking is predicted by a BERT model, fine-tuned on LitBank’s speaker annotations Sims and Bamman (2020). Recently, Vishnubhotla et al. (2022) introduce the largest-to-date corpus of quotation attribution, PDNC, and show a similar accuracy score of around 63% for both BookNLP and the sieve-based model. However, better results were obtained later by fine-tuning BookNLP on PDNC and restricting the predicted coreference chains to resolved characters only Vishnubhotla et al. (2023). SIG Su et al. (2023) adapted quotation attribution for encoder-decoder models, showing that predicting both the speaker and the addressees of a quote with BART reaches state-of-the-art accuracy on PDNC.
Interestingly, reported BookNLP attribution results on PDNC often approach perfect accuracy for explicit quotes, indicating that most of the research efforts should be put in the attribution of non-explicit quotes. Driven by this insight, we exploit explicit quotes uttered by characters to represent them in a latent space, and are the first to show their potential for quotation attribution when combined with contextual representations.
Fictional Character Embeddings.
Understanding fictional characters objectives, personas and style often comes naturally when reading novels. When accessing information from the fictional world and interactions between characters provided throughout the narrative, we infer mental representations of characters that are revised when accessing new information Gernsbacher et al. (1992); Culpeper (1996). Previous computational works on this aspect of fictional character understanding have explored representing fictional characters in a latent space, often targeting a specific aspect, such as persona, style or descriptive elements. These computational representations of characters are often used for further analysis of large literary corpora, trying to understand character distinctiveness between authors Bamman et al. (2014) or narrative comprehension Brahman et al. (2021). Other works focus on using fictional characters embedings to improve character-level tasks such as coreference resolution and character linking Li et al. (2023) or contextual and book-level tasks such as narrative question answering Inoue et al. (2022).
Here, we propose to represent character from what they say and how they say it, building literary embeddings encoding style and topical preferences of characters, derived from models trained on the Authorship Verification (AV) of fictional characters within the same work. Most works exploring the AV of fictional characters focus on drama characters because of the availability of large annotated corpora and aim at capturing syntactic, lexical and phonological variations that occur in their direct speech using statistical measures of distinctiveness Dinu and Uban (2017); Vishnubhotla et al. (2019); Šeļa et al. (2023).
With the recent introduction of PDNC Vishnubhotla et al. (2022), a corpus of 22 novels annotated with quotation attribution3336 additional novels were released a year later, totaling 28 annotated novels., analysis on fictional characters in novels can be done at a larger scale. Michel et al. (2024) provide a comparative study of PDNC’s characters, converting the original PDNC quotation attribution task to the AV of fictional characters and showing that UAR Rivera-Soto et al. (2021) embeddings derived from a small collection of utterances are able to distinguish characters within the same work. However, Reddit – the target domain of UAR – is inherently different from traditional novels that use conventions on grammar, punctuation and format, motivating the need for a variant of UAR tailed for literary texts.
Split Segments Utterances Queries Targets/Query (avg)
Train 1507 263270 5392 5.0
Scene Val 240 50670 1557 8.8
Test 203 41830 1319 8.7
Train 226 449407 4109 90.7
Play Val 30 63934 917 55.1
Test 31 74738 1214 108.5
PDNC Explicit - 6303 562 11.2
Table 1: Summary statistics of the Scene and Play instances of DramaCV by split. A segment in the Scene split is a single scene or act in a play, while a segment in the Play split is the entire play. For each segment, the targets are collections of utterances uttered by all participating characters in the segment. The Explicit split of PDNC Michel et al. (2024) that we use to test generalization is also displayed for comparison.
3 Representing Fictional Characters
In this section, we present our approach to build and evaluate character representations tailored for literary texts, focusing on aspects of style and topical preferences of characters.
To build these representations, we train variants of UAR from scratch on the Authorship Verification (AV) of characters in the literary domain, considering a fictional character as an author in the AV terminology. The AV of fictional characters aims at answering the following problem: given two collections of quotes that have been uttered by characters within the same fictional work, predict whether they were spoken by the same character or not. Although closely linked to Authorship Attribution, we note that this framework is different for the following reasons: 1) we do not have access to a list of canonical characters for every literary work and 2) we compare characters within the same story, implying that each quote is inherently influenced by the author’s style.
UAR adapts SBERT by directly encoding collections of documents, and is trained to distinguish if two collections of documents have been written by the same author or not. Although the Reddit version of UAR appears to perform well on PDNC Michel et al. (2024), which is on-par with its zero-shot domain transfer abilities, we question its suitability for literary analysis and propose to train instead a domain-specific variant of UAR tailored for literary texts.
Only a few datasets containing annotations of the speaker of an utterance of direct speech in full novels exist, as annotating entire works is costly. Instead, prior works have considered using movie Azab2019; Sang et al. (2022); Li et al. (2023) or drama scripts Vishnubhotla et al. (2019); Fischer et al. (2019) as they come naturally annotated with the speaker of each utterance. We follow this approach and propose DramaCV (Drama Character Verification), a dataset of English drama plays focused on the AV of fictional characters. We chose to use drama plays as they are mostly similar to novels in terms of writing conventions, and because we were able to extract a number of characters and utterances an order of magnitude above the available datasets of annotated novels. The rest of this section evaluates our variants of UAR trained on DramaCV on the test set of DramaCV and on fictional characters participating in the 28 PDNC novels.
3.1 DramaCV
We extract 499 plays dating from the 15th to 20th century from Project Gutenberg [^4] https://gutenberg.org/ with the GutenTag software Brooke et al. (2015) that we automatically parsed to create a character list and to attribute each line to its speaker. Then, we construct two instances of DramaCV:
Scene.
We split each play in scenes, a small segment unit of drama that is supposed to contain actions occurring at a specific time and place with the same characters. If a play has no <scene> tag, we instead split it in acts, with the <act> tag. Acts are larger segment units, composed of multiple scenes. For this split, we only consider plays that have at least one of these tags. Although scenes usually contain a few characters, they represent a challenge for CV models as participating characters are akin to discuss a similar topic. Thus, stylistic cues of characters are likely to be useful to distinguish characters within the same scene.
Play.
We do not segment play and use all character lines in a play. Compared to the scene segment, the number of candidate characters is higher, and discussions could include various topics. Hence, topical preferences of characters might be an additional useful feature to distinguish characters.
text
After parsing, a total of 169 plays for the Scene split, and 287 plays for the Play split remained. For both splits, we use 80% of the plays for training, 10% for validation and 10% for test. Following the AV setup, we build queries and targets using collections of character utterances. In this setup, the objective is to recognize if a query and a target have been uttered by the same character or not. We construct queries in the validation and test splits by randomly sampling half of the utterances of a character, and we use the remaining half as the target. The train splits only consider characters that uttered at least 16 lines, such that we can build queries and targets with at least 8 utterances. More details on the construction of queries and targets can be found in Appendix A. Summary statistics for both instances of DramaCV are presented in Table 1.
3.2 Literary UAR
We train UAR from scratch using its public implementation on DramaCV. For both instances, we use the pre-trained all-distilroberta-v1 as the base encoder, and employ a learning rate of 2e-5 for 20 epochs. We set the dimension of the model to $d=512$ and use a maximum sequence length of 64 tokens for all utterances. During training, we build queries and targets by randomly sampling 8 utterances to represent characters. A batch size of 1 play is used when training on the Play split, and 8 scenes are used when training on the Scene split. During inference, we do not restrict the number of utterances in a collection to reflect situations where we have access to only less than 8 utterances, or to more than 8 utterances.
A supervised contrastive objective Khosla et al. (2020) is used where in-batch negatives considered are characters within the same segment unit. We motivate this choice for each dataset instance:
Scene.  In a scene, characters are likely to discuss similar topics. By restricting the objective to characters within a scene, we force the model to rely on stylistic cues to distinguish lines of characters.
Play. We want to avoid capturing the writing style of authors to distinguish between collections of character lines. Thus, we restrict the set of in-batch negatives to characters within the same work.
We compare our variants of UAR, $\text{UAR}_{Scene}$ [^5] https://huggingface.co/gasmichel/UAR_Scene and $\text{UAR}_{Play}$ [^6] https://huggingface.co/gasmichel/UAR_Play against their Reddit counterpart, $\text{UAR}_{Reddit}$ and SBERT. Following previous work Aggazzotti et al. (2023); Michel et al. (2024), we create SBERT character embeddings by encoding independently each utterance in a feature $d$-dimensional vector [^7] We use all-mpnet-base-v2., followed by a coordinate-wise mean of all utterance vectors. Instead, UAR directly encodes a collection of utterances into a $d$-dimensional vector.
Given a collection of a character utterances (query) and collections of all characters occurring in the same segment (targets, that include a target uttered by the same character as the query), we evaluate the ability of these models to yield a higher cosine-similarity to the target uttered by the same character than to targets uttered by other characters using Area Under the Receiver Operating Characteristic curve (AUC). We use this evaluation setup on fictional characters for both test-sets of DramaCV and for the 28 novels of PDNC on the Explicit split of Michel et al. (2024). More details on the evaluation can be found in Appendix A.
DramaCV PDNC - Exp
Scenes Plays CC CQ
SBERT 72.0 (5) 78.8 (10) 63.8 (16) 54.4 (5)
$\text{UAR}_{Reddit}$ 68.6 (7) 74.5 (11) 81.1 (9) 54.1 (4)
$\text{UAR}_{Scene}$ 82.3 (6) - 81.3 (10) 58.9 (4)
$\text{UAR}_{Play}$ 84.2 (9) 86.5 (8) 58.5 (5)
Table 2: Average AUC (%) on test plays in DramaCV and on novels in the Explicit split of PDNC with standard deviations between parentheses.
3.3 Authorship Verification Results
We report AUC for both test splits of DramaCV and the Explicit split of PDNC in Table 2. Interestingly, $\text{UAR}_{Reddit}$ performs worse on both splits compared to SBERT and in-domain UARs. $\text{UAR}_{Scene}$ increases AUC on its training domain up to 10 points compared to SBERT, indicating that the model is able to capture stylistic cues specific to drama that help at distinguishing utterances of characters discussing similar topics. The increase of $\text{UAR}_{Play}$ is slightly lower, but the high AUC indicates that it succeeds at distinguishing characters within the same play. The high standard deviation across plays suggests that task difficulty is not distributed uniformly among these literary works. Among best-performing plays, we found works by canonical authors such as Shakespeare’s Hamlet, Shaw’s John Bull’s Other Island or Ben Jonson’s The Alchemist (see complete list in Appendix B).
Looking at generalization results of PDNC, $\text{UAR}_{Reddit}$ performs better than SBERT, and almost equals $\text{UAR}_{Scene}$ in the Character-Character setup (CC), which compares similarity of character representations. We find significant improvement with $\text{UAR}_{Play}$ over the baselines, suggesting that training UAR on drama allows better domain adaptation to literary novels. We also remark that this adaptation is data-efficient, as training $\text{UAR}_{Scene}$ with around 5000 drama characters is enough to match the performance of $\text{UAR}_{Reddit}$ that has been trained on millions of different users.
The advantages of domain adaptation can also be seen in the Character-Quote (CQ) evaluation setup that aims at attributing utterances to their speaker by comparing the similarity between character and quote embeddings computed with the same model. Here, the two variants of UAR trained on drama show a significant increase in AUC over the baselines, indicating that character representations built with our proposed UARs are more similar to embeddings of their other quotes. However, AUC scores for the CQ experiment are low, showing that the task of attributing quotes requires more than only comparing similarities of character and quote embeddings.
Figure 2: Detailed process of BookNLP’s quotation attribution system. We inject extra-information in the form of character embeddings computed from explicit quotes $v_{\text{entity}(m)}$ in the fourth step, and quote embeddings $u_{q}$ in the second step.
4 Quotation attribution
Motivated by the ability of UAR to encode character utterances in a representation space that contains useful, though insufficient, information for quotation attribution, we evaluate in the next section the impact of combining these global character embeddings with BookNLP’s contextual information. BookNLP is a Natural Language Processing pipeline dedicated to English books, popular among practitioners of Digital Humanities.
4.1 BookNLP
BookNLP’s pipeline includes, among others, Part-of-speech tagging, Named Entity Recognition, Coreference Resolution and quotation attribution. All components of the pipeline have been trained on LitBank Bamman et al. (2019, 2020); Sims and Bamman (2020), a corpus containing 100 chapters from public domain books. Since its quotation attribution model is not part of a publication, we will review its details in the following paragraphs.
BookNLP carries out quotation attribution in four independent steps: 1) tokenization and named entity recognition, 2) character name clustering, 3) pronominal coreference resolution, and 4) quote-mention linking. The first three steps are dedicated to extract named, nominal and pronominal mentions along with the predicted character they refer to. Using extracted mentions in the surrounding context of a quote, the final step, quote-to-mention linking, aims at predicting which mention is the one that refers to the true speaker of the quote. In other words, it finds the information given by the narrator about who is speaking among the narrative parts in the vicinity of quotes. It is achieved by a fine-tuned BERT model Devlin et al. (2019). We give further details of the quote-mention linking step below.
Overall Non-Explicit Explicit Anaphoric Implicit
SIG $72^{+}$ $70.0^{+}$
ChatGPT $71^{+}$ $70.0^{+}$
Unanswerable (%) 14.1 (4.4) 20.9 (6.7) 0 24.5 (8.1) 19.8 (8.6)
[…]
+ SBERT $79.2^{\star\star}$ (3.8) $70.0^{\star\phantom{\star}}$ (4.5) 98.3 (1.1) $70.8^{\phantom{\star\star}}$ (7.3) $68.1^{\star}$ (5.5)
+ $\text{UAR}_{Reddit}$ $80.1^{\star\star}$ (3.9) $71.1^{\star\star}$ (4.7) 98.7 (1.1) $71.5^{\phantom{\star\star}}$ (7.9) $\mathbf{69.6}^{\star}$ (5.3)
+ $\text{UAR}_{Play}$ $80.0^{\star\phantom{\star}}$ (3.8) $70.9^{\star\phantom{\star}}$ (4.4) 98.7 (1.0) $71.1^{\phantom{\star\star}}$ (7.9) $69.5^{\star}$ (5.7)
+ $\text{UAR}_{Scene}$ $\mathbf{80.2}^{\star\phantom{\star}}$ (3.8) $\mathbf{71.2}^{\star\phantom{\star}}$ (4.3) 98.7 (1.0) $\mathbf{71.7}^{\star\star}$ (8.1) $\mathbf{69.6}^{\star}$ (5.4)
[…]
Let $D=(t_{1},\dots t_{n})$ be a tokenized document and $q=(t_{i},\dots,t_{j})$ be a character utterance in this document, starting at token $i$ and ending at token $j$. A contextual segment is computed for each quote, $c_{q}=[\,c_{q}^{left}\;|\,q\,|\;c_{q}^{right}\,]$, where $c_{q}^{left}$ and $c_{q}^{right}$ are left and right contextual information of length $w$ tokens and [|][\,|\,][ | ] denotes concatenation. Additionally, the full quote $q$ is replaced by a special token [QUOTE] and all other quotes occurring in the contextual segment are replaced by a special token [ALTQUOTE], such that the contextual segment mainly reflect narrative parts. Each segment $c_{q}$, containing masked quotes and narrative parts, is fed to BERT, producing a contextual representation $\mathbf{H}_{c_{q}}$. Then an unary compatibility score between the (masked) quote token $q$ and a mention-span $m=(t_{m_{s}},\dots,t_{m_{e}})$ occurring in the segment is calculated as follows:
$h_{m}$ $=\frac{1}{|m|}\sum\limits_{i=m_{s}}^{m_{e}}\mathbf{H}_{c_{q}}[i]$ (1)
$s(q,m)$ $=\phi([\mathbf{H}_{c_{q}}[q]\;|\;h_{m}])$ (2)
where $\mathbf{H}_{c_{q}}[i]$ is the BERT representation of token $i$ while $|m|$ is the token length of mention $m$ and $\phi$ is a feed-forward neural network with one hidden layer and a tanh activation. Given a set of candidates $Y(q)=(m_{1},\dots,m_{k})$, the model is trained to optimize the log-likelihood of mentions referring to the true speaker of $q$. Figure 2 summarizes this process.
The default window size is $w=50$ tokens, and the hidden layer dimension is $100$. During training and inference, only a maximum of $k=10$ mention candidates are considered.
In our experiments, we modify these default parameters by increasing the narrative window size to $w=100$ tokens, and using a relu activation and a hidden dimension of $d=512$ in the feed-forward neural network. We also replace BERT by SpanBERT-large Joshi et al. (2020), computing mention representations by concatenating SpanBERT embeddings of the first and last tokens. At inference, we do not truncate the number of candidates and instead use all candidate mentions available in the context window.
4.2 Integrating Character Embeddings
Let $E(c)=(q_{i},\dots,q_{j})$ be the collection of all explicit quotes uttered by character $c$. We embed this collection of quotes using one of the model described in Section 3.2, yielding a vector representation $v_{c}$ for each character. If a character has not been explicitly quoted, we set its vector representation to zeros. Additionally, we embed each target quote $q$ independently with the same model to get a vector representation $u_{q}$. We modify Equation 2 to combine BookNLP’s contextual representation with these character and quote representations:
$s(q,m)=\phi([\mathbf{H}_{c_{q}}[q]\;|\;h_{m}\;|\;v_{\text{entity}(m)}\;|\;u_{q}])$ (3)
where $\text{entity}(m)$ is the character predicted by the coreference resolution step for mention $m$.
Equation 3 allows BookNLP to rely on external information by injecting global character information and compatibility between quote $q$ and candidate character $c$.
4.3 Experimental Setup
Following Su et al. (2023), we use the first version of PDNC, which contains 22 annotated novels. We focus on major and intermediate characters, who are characters that uttered at least 10 quotes in the novel. This allows to avoid the long tail of minor characters and to compare our results to the current state-of-the-art, SIG and ChatGPT (gpt-3.5-turbo-0613).
We base our implementation on BookNLP+ Vishnubhotla et al. (2023), that restricts coreference predictions to chains that can be successfully resolved to a character entity from the annotated PDNC character lists. We follow the same evaluation protocol, using their provided cross-validation split at the novel level, reporting the average accuracy over each split.
[…]
4.4 Results
[…]
Combining BookNLP+ with SBERT character embeddings matches the accuracy of SIG and ChatGPT on non-explicit quotes. The gains are larger for all variants of UAR, reaching 71.2% accuracy on non-explicit quotes with our proposed $\text{UAR}_{Scene}$. As expected, we found that these gains are largely due to the improved attribution performance of implicit quotes, increasing the accuracy by more than 3 points with all UAR models.
Figure 3: Accuracy results on non-explicit quotes averaged over subsets of novels that have a low percentage of unanswerable quotes, with standard deviations.
Interestingly, although our proposed variants of UAR tailored for literary texts are better at the Authorship Verification of fictional characters, we only see a small – though statistically significant – increase between $\text{UAR}_{Scene}$ and its Reddit counterpart on non-explicit attribution accuracy. The reasons could be twofold. First, BookNLP’s system only focuses on finding the right narrative information that helps identifying the speaker of a quote. However, our embeddings mostly reveal stylistic cues of characters, which are not encoded by BookNLP. Therefore, our implementation provides external character information that might not align perfectly with BookNLP’s contextual decisions, limiting their potential combination. Second, we add character information based on noisy coreference chains, which might hinder the model’s performance. We believe further research should be conducted into finding better combinations between character embeddings and BookNLP.
The high standard deviations across splits reveals that the task is harder for some novels than others. Looking at the percentage of unanswerable quotes (quotes that do not have a candidate mention referring to their speaker in their context window) provides a reason for such high variations: it ranges from 0 to 56% per novel. Figure 3 shows the attribution accuracy of non-explicit quotes when only considering novels that have a low percentage of unanswerable quotes. Only 4 novels have a percentage lower than 5%, but 17 out of the 22 have a percentage lower than 20%. As expected, we see higher performance for all models, and the gain from UAR character embeddings is more clear.
$\Delta$ Non-Exp $\Delta$ Ana $\Delta$ Imp
SBERT $-0.0\phantom{{}^{\star\star}}$ $-0.0\phantom{{}^{\star}}$ $-0.1\phantom{{}^{\star}}$
$\text{UAR}_{Reddit}$ $-0.1\phantom{{}^{\star\star}}$ $-0.0\phantom{{}^{\star}}$ $-0.2\phantom{{}^{\star}}$
$\text{UAR}_{Play}$ $-0.2^{\star\star}$ $-0.1\phantom{{}^{\star}}$ $-0.3^{\star}$
$\text{UAR}_{Scene}$ $+0.1\phantom{{}^{\star\star}}$ $+0.1\phantom{{}^{\star}}$ $+0.1\phantom{{}^{\star}}$
[…]
Let us note that $\text{UAR}_{Scene}$ is consistently better than $\text{UAR}_{Play}$, though the latter is better at distinguishing PDNC characters in the CC experiment. Restricting UAR to encode stylistic information by using collections of quotes with similar topic seems to generalize better to quotation attribution.
[…]
Finally, we present quotation attribution results for 6 annotated novels from the second release of PDNC in Table 5. We use the model trained on the first cross-validation split to predict attributions in each of these novels. The impact of using literary character embeddings is more clear: both of our proposed variants of UAR show increased attribution accuracy against their Reddit counterpart. Besides, the increase in accuracy over non-explicit quotes is again mostly attributed to an increase in implicit speaker identification.
5 Conclusion
In this work, we evaluated fictional character embeddings constructed with the explicit quotes they utter, showing that they improve quotation attribution accuracy of non-explicit quotes on PDNC when used in combination with BookNLP’s contextual representations. We created DramaCV, a corpus of English drama plays focused on the Authorship Verification of fictional characters and trained two variants of UAR on this dataset. Our proposed models show better performance when distinguishing fictional characters in both DramaCV and PDNC. Character embeddings derived from these models are further used for quotation attribution on PDNC, reaching state-of-the-art accuracy when combined with BookNLP+. The impact of using literary character embeddings against embeddings derived from a model trained on Reddit is small, though statistically significant, suggesting superiority of in-domain character representations for quotation attribution.
We proposed a straightforward way to combine fictional character embeddings with contextual representations, using BookNLP as an example. Our approach can be adapted to other models: cross-attention over character embeddings could be used for encoder-decoder models, while modality alignment Zhang et al. (2023) is a solution for decoder-only models. We leave potential improvements to the combination of fictional character embeddings and contextual representations to future work.
Non-Exp Ana Imp
Unanswerable (%) 20 21 18
[…]
+ SBERT $69.3^{\phantom{\star\star}}$ $71.6^{\phantom{\star\star}}$ $68.0^{\phantom{\star}}$
+ $\text{UAR}_{Reddit}$ $70.5^{\star\star}$ $72.4^{\phantom{\star\star}}$ $69.4^{\star}$
+ $\text{UAR}_{Play}$ $71.1^{\star\phantom{\star}}$ $72.8^{\star\star}$ $\mathbf{70.1}^{\star}$
+ $\text{UAR}_{Scene}$ $\mathbf{71.2}^{\star\phantom{\star}}$ $\mathbf{73.1}^{\phantom{\star\star}}$ $70.0^{\star}$
[…]
We found that BookNLP’s attribution accuracy is not distributed uniformly across novels, mainly caused by noisy coreference resolution, resulting in many quotes not having contextual candidate mentions that refer to their true speaker. We believe that including our character embeddings in the coreference resolution model as in Li et al. (2023) might also improve its performance and thus increase overall attribution accuracy. Another potential solution could include unmasking the closest quotes in the vicinity of a target quote that are likely to contain addressee information. We leave better combination of fictional character embeddings with BookNLP for future exploration.
6 Limitations
In this work, we focused on integrating fictional character embeddings in encoder-only models. We have not tested with other types of models such as encoder-decoder and decoder only, but we believe that integrating our embeddings in those frameworks is possible and is likely to improve their performance on the quotation attribution task. Our first set of experiments corroborate this assumption, on the usefulness of our proposed character embeddings in distinguishing fictional characters.
We also proposed models that can be used to create character representations. These models are trained on a corpus of drama plays, which spans almost 400 years of work. It is unclear how the variation of writing styles that have occurred throughout these centuries affects the information contained in these embeddings, and how they can generalize to new literary styles. This remark also extends to our results on quotation attribution. We presented results on a corpus containing novels from the 19th to early 20th century ; which is closer in terms of writing style to DramaCV than newer novels published in the 21th century. It is thus unclear how our results generalize to these newer novels, but we strongly believe that accurate character embeddings improve quotation attribution regardless of the writing style of the author.
We only focused on representations encoding stylistic cues and topical preferences and showed that they improve quotation attribution. However, this type of character embedding might be incomplete and other aspects of characters can be used in addition to our proposed embeddings (persona, gender, objective, or social status). This might allow to build accurate representations of characters that are likely to improve even more quotation attribution accuracy.
7 Acknowledgements
This work was partly performed using HPC resources from GENCI-IDRIS and Grid500.
Appendix A Details on the Construction and Evaluation of DramaCV
Each play collected with GutenTag is encoded in a ‘XML’ file, containing various tags such as the <speaker> tag that indicates the speaker of a line. We parse each XML file, removing plays written by non English-speaking authors, and attribute each line to its speaker using the <speaker> tag.
For the Scene split, we automatically check if a play contains a <scene> or <act> tag. We segment plays by default with the <scene> tag, and use the <act> tag if it does not contain a scene tag. We removed plays that did not contain any of these tags. For the Play split, we kept all plays written by English-speaking authors.
In the Authorship Verification terminology, queries and targets are collections of documents written by either the same or different authors. Given a query written by author A and multiple targets written by either the same or different authors, the objective of AV is to recognize all targets written by author A. Details of queries and target construction and our evaluation setup are provided below:
Queries/Targets.
Given characters $(c_{1},\dots,c_{n})$ occurring in a single segment (a scene/act for the Scene split, or the entire play for the Play split), we randomly sample half of their utterances to construct queries $(q_{c_{1}},\dots,q_{c_{n}})$ and use the remaining half to construct targets $(t_{c_{1}},\dots,t_{c_{n}})$. For the train split, we instead randomly sample 8 utterances to build queries and 8 different utterances to build targets.
DramaCV Evaluation.
Using any model $\mathcal{M}$ that is able to encode collections of documents in a feature vector, we encode each query and target independently. Then, we compare the cosine similarity of query vectors with targets: $similarity(\mathcal{M}(q),\mathcal{M}(t))$ $\forall q\in q_{c_{1}},\dots,q_{c_{n}}$ and $\forall t\in t_{c_{1}},\dots,t_{c_{n}}$. For a query $q_{c}$ uttered by character $c$, we calculate an AUC score over the resulting cosine similarities with targets, where a label $1$ is assigned to $t_{c}$ and 00 to other targets. This AUC represents the ability of $\mathcal{M}$ to assign an higher cosine similarities between queries and targets written by the same author. We average this AUC score over each query in a single segment to obtain an AUC score for this particular segment.
PDNC Evaluation.
We use the evaluation setup of Michel et al. (2024) to evaluate our models on the characters of PDNC. This setup proposes two types of evaluation: Character-Character (CC) and Character-Quotes (CQ).
The CC evaluation is similar to the one described above, but the queries are constructed using only the explicit quotes of a character in a single chapter, and targets are constructed using every other utterance in the other chapters. The AUC computed scores how a model is able to assign a higher cosine-similarity between a collection of a few explicit quotes of character, and its other utterances in the novel.
In the CQ setup, queries are similar to CC, but a target is defined as a single utterance in the held-out chapters, without access to contextual information. Therefore, the AUC computed is scoring how similar the collection of a few explicit quotes of character is to its other individual quotes. This task is inherently harder than the CC setup, as an individual quote might be very different from other quotes (e.g “Hey, how are you Lizzy?”).
Appendix B Performance on all Test Plays
We report performance of $\text{UAR}_{Scene}$ and $\text{UAR}_{Play}$ on their respective test split in Table 6 and Table 7.
Appendix C Computing Information
[…]
Play Author AUC
Justice John Galsworthy 95.6
The Alchemist Ben Jonson 91.8
Hamlet William Shakespeare 86.5
Plays : Third Series John Galsworthy 86.3
The works of John Dryden John Dryden 85.2
The Life of Timon of Athens William Shakespeare 84.2
King Henry the Fifth William Shakespeare 83.7
Night Must Fall : a Play in Three Acts Emlyn Williams 83.7
The Spanish Tragedy Thomas Kyd 83.2
The works of John Dryden John Dryden 83.0
The Scarlet Stigma: A Drama in Four Acts Nathaniel Hawthorne 82.4
King John William Shakespeare 81.8
The Works of Charles and Mary Lamb Mary Lamb 80.8
The Scornful Lady John Fletcher 80.8
Married Life A Comedy, in Three Acts John Baldwin Buckstone 80.0
The Flutter of the Goldleaf; and Other Plays Frederick Peterson 77.5
Single Life A Comedy, in Three Acts John Baldwin Buckstone 75.9
The Countess Cathleen W. B. (William Butler) Yeats 75.6
De Turkey and De Law: A Comedy in Three Acts Zora Neale Hurston 66.4
Table 6: AUC per novel on the Scene test split of DramaCV.
Play Author AUC
Bride Roses William Dean Howells 100.0
Fanny’s First Play Bernard Shaw 100.0
Five Little Plays Alfred Sutro 98.6
John Bull’s Other Island Bernard Shaw 95.4
Mrs. Dot: A Farce William Somerset Maugham 94.5
The Earl of Essex: A Tragedy, in Five Acts Henry Jones 94.4
The Way of the World William Congreve 92.7
Plays : Fifth Series John Galsworthy 92.5
Representative Plays by American Dramatists: 1856-1911: The New York Idea Langdon Elwyn Mitchell 91.9
Plays: Lady Frederick, The Explorer, A Man of Honour William Somerset Maugham 90.3
Great Catherine Bernard Shaw 89.3
Hypolympia Or, The Gods in the Island, an Ironic Fantasy Edmund Gosse 87.1
The works of John Dryden John Dryden 87.0
The Grecian Daughter Arthur Murphy 85.5
Dramatic Technique George Pierce Baker 81.6
Lyre and Lancet: A Story in Scenes F. Anstey 80.2
Beaumont and Fletcher’s Works, Vol. 5 Francis Beaumont 79.0
King Henry VI, Part 3 William Shakespeare 78.1
Plays and Lyrics Cale Young Rice 78.0
The Easiest Way Representative Plays by American Dramatists: 1856-1911 Eugene Walter 77.6
Charles Di Tocca: A Tragedy Cale Young Rice 77.3
St. Patrick’s day, or, the scheming lieutenant : a farce in one act Richard Brinsley Sheridan 76.7
Deirdre of the Sorrows John Millington Synge 76.3
Ambrose Gwinett or, a sea-side story : a melo-drama, in three acts Douglas William Jerrold 75.8
Henry V William Shakespeare 75.3
Box and Cox: A Romance of Real Life in One Act. John Maddison Morton 75.0
The Recruiting Officer George Farquhar 74.7
Yolanda of Cyprus Cale Young Rice 74.3
Table 7: AUC per novel on the Play test split of DramaCV.